OpenStudy (anonymous):

Last factor question! Factor completely. 3x3 – 48x

OpenStudy (anonymous):

i thnk its 3x(x – 16)

OpenStudy (anonymous):

It should be x^2 - 16 not x - 16

OpenStudy (anonymous):

3x^3 - 48x = 3x(x^2 - 16) = 3x(x - 4)(x + 4) One of the basic algebraic identity a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

Oh well its out of these answers A. 3x(x – 4)(x + 4) B. 3x(x – 4)(x – 4) C. 3(x + 4)(x – 4) D. 3x(x – 16)

OpenStudy (anonymous):

I see now thanks